OpenStudy (anonymous):

What is the square root of 180? (Simplify) I'm new to this subject and if possible could someone explain to me this process

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

sqrt(180)...to simplify ...you need to find, if possible, a common factor that divides 180 and is a perfect square such as 4,9,16,etc..

OpenStudy (anonymous):

Ok Thank you

OpenStudy (anonymous):

You can see that 180=18*10=2*4*9

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Do not forget the medal please :D